# Freeze and Conquer: Reusable Ansatz for Solving the Traveling Salesman Problem

### v1

@@ 3.1 Ansatz Optimization
Selecting an appropriate $\mathcal{A}$ size is challenging because excessively large circuits incur high computational costs, while undersized ones produce suboptimal results. To tackle this issue, we partition our $\mathcal{A}$ into 5 distinct blocks, each of which can be either a Rotation block or an Entanglement block. Rotation blocks consist of rotation gates ($R_{x}$, $R_{y}$, $R_{z}$), whereas Entanglement blocks are constructed using $CX$ gates (see [21]). The first block of $\mathcal{A}$ is always a rotation block, followed by an entanglement block, and this alternating pattern continues until the final block, which is also a rotation block. Within each rotation block, every gate has an independent parameter $\Theta_{i}[j]$, where $i$ denotes the block index and $j$ represents the qubit index.
@@
This $\mathcal{A}$ is represented as a quantum circuit in Figure 1. The first block consists of $R_{z}$ gates, one for each qubit, forming a rotation layer. This is followed by an entanglement block composed of linearly arranged $CX$ gates. The pattern alternates between rotation and entanglement blocks, ending with a final rotation block of $R_{y}$ gates.
@@ 4.1 Problem setup
We generate TSP instances using a controlled procedure, where each instance is represented by a symmetric $n\times n$ distance matrix, with $n$ denoting the number of cities. Distances between cities are randomly sampled as integers within a predefined range (e.g., $1$ to $100$), while self-distances are set to 0 to ensure validity. This generation method enables the efficient construction of instances for testing and comparing algorithms across varying levels of complexity and problem size.
@@ 4.3.1 Convergence to the optimal solution
Concretely, each qubit is first rotated around the $z$ axis ($R_{z}$). These single qubit phases are then mixed by a fully connected $CX$ block, which connects each qubit with all others. A second layer of $R_{x}$ refines the phases introduced so far, after which a $CX$ scheme provides linear entanglement. The circuit closes with the $R_{z}$ layer.
@@ 4.3.2 Reusing
The deterioration becomes evident as it grows to seven cities. As Tab. 2 shows, with a very permissive success threshold, $\tau=0.20$, the algorithm manages to sample the optimal tour in only $3$ of the $10$ test instances. Increasing the threshold to $\tau=0.40$ reduces the successes to $2$, and the count remains stuck at $2$ for all subsequent thresholds. In other words, in $80\%$ times of the seven-city instances the probability of extracting the optimum remains below $0.60\%$, and in $60\%$ of the cases it never exceeds $0.20\%$. This sharp decline confirms that the current method does not efficiently generalize the larger search space created.
@@ 4 Cities 5 Cities 6 Cities 7 Cities
For the most challenging instances at $7$ city (orange box in Fig. 4) performance plummets. The median accuracy drops to $\approx 0.20$, with first quartile $Q_{1}\!\simeq\!0.02$ and third quartile $Q_{3}\!\simeq\!0.25$, so that the inter-quartile range holds of $0.23\%$. Both “whiskers” remain anchored at the low end of the axis (whisker lower than 0, upper $0.25$) and only two high outliers appear, around $\approx 0.90$. In probabilistic terms, this implies that at least $75\%$ of the runs sample the optimal tour with probability less than $0.25$, indicating that the generalization ability of the circuit deteriorates almost completely when the problem size reaches seven cities.

### v2

@@ 3.1 Ansatz Optimization
Selecting an appropriate $\mathcal{A}$ size is challenging because excessively large circuits incur high computational costs, while undersized ones produce suboptimal results. To tackle this issue, we partition our $\mathcal{A}$ into 5 distinct blocks, each of which can be either a Rotation block or an Entanglement block. Rotation blocks consist of rotation gates ($R_{x}$, $R_{y}$, $R_{z}$), whereas Entanglement blocks are constructed using CXCX gates (see [21]). The first block of $\mathcal{A}$ is always a rotation block, followed by an entanglement block, and this alternating pattern continues until the final block, which is also a rotation block. Within each rotation block, every gate has an independent parameter $\Theta_{i}[j]$, where $i$ denotes the block index and $j$ represents the qubit index.
@@
This $\mathcal{A}$ is represented as a quantum circuit in Figure 1. The first block consists of $R_{z}$ gates, one for each qubit, forming a rotation layer. This is followed by an entanglement block composed of linearly arranged CXCX gates. The pattern alternates between rotation and entanglement blocks, ending with a final rotation block of $R_{y}$ gates.
@@ 4.1 Problem setup
We generate TSP instances using a controlled procedure, where each instance is represented by a symmetric $n\times n$ distance matrix, with $n$ denoting the number of cities. Distances between cities are randomly sampled as integers within a predefined range (e.g., 11 to $100$), while self-distances are set to 0 to ensure validity. This generation method enables the efficient construction of instances for testing and comparing algorithms across varying levels of complexity and problem size.
@@ 4.3.1 Convergence to the optimal solution
Concretely, each qubit is first rotated around the $z$ axis ($R_{z}$). These single qubit phases are then mixed by a fully connected CXCX block, which connects each qubit with all others. A second layer of $R_{x}$ refines the phases introduced so far, after which a CXCX scheme provides linear entanglement. The circuit closes with the $R_{z}$ layer.
@@ 4.3.2 Reusing
The deterioration becomes evident as it grows to seven cities. As Tab. 2 shows, with a very permissive success threshold, $\tau=0.20$, the algorithm manages to sample the optimal tour in only 33 of the $10$ test instances. Increasing the threshold to $\tau=0.40$ reduces the successes to 22, and the count remains stuck at 22 for all subsequent thresholds. In other words, in $80\%$ times of the seven-city instances the probability of extracting the optimum remains below $0.60\%$, and in $60\%$ of the cases it never exceeds $0.20\%$. This sharp decline confirms that the current method does not efficiently generalize the larger search space created.
@@ 4 Cities 5 Cities 6 Cities 7 Cities
For the most challenging instances at 77 city (orange box in Fig. 4) performance plummets. The median accuracy drops to $\approx 0.20$, with first quartile $Q_{1}\!\simeq\!0.02$ and third quartile $Q_{3}\!\simeq\!0.25$, so that the inter-quartile range holds of $0.23\%$. Both “whiskers” remain anchored at the low end of the axis (whisker lower than 0, upper $0.25$) and only two high outliers appear, around $\approx 0.90$. In probabilistic terms, this implies that at least $75\%$ of the runs sample the optimal tour with probability less than $0.25$, indicating that the generalization ability of the circuit deteriorates almost completely when the problem size reaches seven cities.